# MLVU: A Comprehensive Benchmark for Multi-Task Long Video Understanding

### v2

@@ 3.2.2 Single-Detail LVU
Needle Question-Answering (NQA). Needle-In-the-Haystack-Search (NIHS) is a popular evaluation task for long-context LLM [26]. Taking the inspiration from NIHS, we create Needle Question-Answering (NQA), shown as Figure 2 (d). In this task, the MLLM is required to answer a question related to a specific segment (referred as needle) within a long video (referred as background video). The needles are short video clips sampled from WebVid [5], while the background videos are sampled from our ULVC. The needle is randomly inserted into the background video, where a question-answer pair is annotated. By incorporating necessary details, the question can always correspond to the needle without ambiguity. During evaluation, the MLLM needs to infer the location of the needle based on the details provided in the question, and solve the problem on top of the needle’s information. The NQA task is structured as multiple-choice, whose performance is measured by accuracy.
Ego Reasoning (ER). Ego-centric videos capture a series of consecutive actions from a first-person perspective. The MLLM needs to reason for a question about a specific behavior in the video, e.g., predicting for the event which is correlated or satisfies a certain causal relationship with the behavior (Figure 2 e). Both videos and QA annotations are collected from the NLQ task of Ego4D [11]. The ER task is structured as multiple-choice, with a total of 352 questions created for this task.
Plot Question-Answering (PQA). In this task, the MLLM needs to reason for questions about a plot in a narrative video, shown as Figure 2 (f). The video is sampled from the movies, TV series, and animated series in our ULVC. There are 539 question-answer pairs created by manual annotation. During annotation, the human annotators are asked to only provide necessary details about the plot but not to suggest any objective hints, e.g., the two characters in the example video are referred as cat and mouse, rather than Tom and Jerry. Therefore, it can prevent the question from being short-cut by the MLLM’s common-sense knowledge (more details about PQA can be found in the Appendix E.6).
Sub-Scene Captioning (SSC). In this task, the MLLM needs to generate the caption for a sub-scene in a long video. The long videos in SSC are sampled from the Movie101 dataset [52], while the questions and answers are manually annotated. During annotation, the human annotator is asked to provide a detailed description for the sub-scene as the ground-truth answer. Besides, they need to offer necessary clues in their questions such that the referred sub-scenes can be identified without ambiguity. During evaluation, we employ GPT-4 [1] to measure the quality of caption in comparison with the ground-truth. Details about annotation and evaluation are presented in Appendix E.7 and F.3.
@@ 3.2.3 Multi-Detail LVU
Action Order (AO). In this task, the MLLM needs to predict the right order for a sequence of actions (Figure 2 h). The actions are presented by short video clips, called probes. The probes are formulated in two different ways. One is made up of clips from the Kinetics dataset [14], where each clip represents a distinct action. The other one is from the consecutive clips of an action in the ActivityNet-Caption dataset [16]. The probes are inserted into a long background video, which is sampled from ULVC. There are 259 AO questions in total. The task is structured as a multiple-choice prblem, where the right order is selected from the misleading options provided by the annotator.
Action Count (AC). This task requires the MLLM to count the occurrences of an action within a long video (Figure 2 i). Each action corresponds to multiple short probe clips sampled from the Kinetics dataset [14]. The probes of an action are inserted into a long background video sampled from ULVC. We also perform manual examination to ensure that the inserted action does not exist in the original background video. A total of 206 evaluation instances have been created. The AC task is structured as a multiple-choice problem, with performance measured by accuracy.
@@ 4.1 Settings
We perform a comprehensive investigation of 20 MLLMs based on MLVU, both open-source and proprietary. The experimental MLLMs can be partitioned into three categories. 1. Image MLLMs, which are primarily fine-tuned by image-related instructions. We consider the following models for this category: Otter-I [18], LLaVA-1.6 [24], InternVL-1.5 [7], Claude3-Opus [2], Qwen-VL-Max [4], and GPT-4 Turbo [1]). 2. Short Video MLLMs, which are fine-tuned by short-video related instructions. This category includes: Otter-V [18], mPlug-Owl-V [49], Video LLaMA-2 [53], Video ChatGPT [29], VideoChat [20], VideoChat2 [21], and Video-LLaVA [23]. 3. Long Video MLLMs, which are optimized for their long-video understanding capability. This category includes: MiniGPT4-Video [3], LLaMA-VID [22], Movie-LLM [37], MA-LMM [12], MovieChat [36], TimeChat [35], and GPT-4o [33]. For Image MLLMs, we leverage their multi-image inference capabilities to process segmented frames from original videos. In the case of Video MLLMs, we employ either a uniform sampling strategy or a frame rate sampling strategy for video processing. All models are evaluated based on their official implementations or available APIs, where the evaluation is conducted in a zero-shot manner. More details about the evaluation are provided in the Appendix F.
4.2 Overall Performance
@@
The overall evaluation results for all investigated MLLMs are shown in Table 1. The individual performance is reported for each individual task; meanwhile, the average performances are reported for the multiple-choice (M-Avg) and generation tasks (G-Avg), respectively. We can obtain two primary observations from the demonstrated results.
On one hand, the recently released GPT-4o [33] achieves an overwhelming advantage in our benchmark. It outperforms the rest of methods by a big margin in terms of the average performances, with an M-Avg of 64.6% (within 0-100%) and a G-Avg of 5.80 (within 0.0-10.0). Besides, it also maintains the leading position in every individual task. As for the open-source models, InternVL-1.5 [7] achieves the highest average performance in multiple-choice tasks, whose M-Avg scores reaches 50.4%. Such a performance is impressive knowing that it even slightly goes beyond GPT-4 Turbo [1]. Meanwhile, LLaMA-VID [22] presents the top performance for the generation tasks, whose G-Avg score reaches 4.22. However, such a score is significantly lagging behind GPT-4 Turbo and GPT-4o.
@@
On the other hand, although GPT-4o achieves a huge advantage in comparison with other methods, it actually struggles to handle most of the tasks in the benchmark. For example, it only ends up with 64.8% in dealing with the needle question-answering (NQA) task. Whereas its analogue tasks in the text domain, e.g., NIHS (Needle-In-the-HayStack-Search) and Passkey Retrieval, can be effectively conquered by many of the existing long LLMs [10; 54]. At the same time, it exhibits even less reliability given tasks like ego-reasoning (ER), action ordering (AO), and action count (AC), while other baseline methods produce even worse performances in these scenarios. The above observations indicate that long-video understanding remains a tough challenge for today’s MLLMs.
@@
In addition to the primary conclusions from the overall performances, we can also make the following interesting observations about the individual tasks. First of all, the multiple-choice holistic tasks, i.e., topic retrieval (TR) and anomaly recognition (AR), present much higher differentiation than other tasks. Proprietary MLLMs, like GPT-4o, GPT-4-turbo, and superior open-source models, like InternVL-1.5, can accurately solve such problems; meanwhile, many other popular MLLMs still fail to generate meaningful performances. Knowing that the two tasks only require an overall understanding of the long videos, they can serve as a preliminary indicator of MLLMs’ LVU ability.
@@
Besides, it’s hard to deal with tasks which need nuanced understanding of multiple details. Although several MLLMs can handle single-detail LVU tasks to some extent, their performances suffer from catastrophic degradation when addressing multi-detail LVU tasks. Most methods, except for GPT-4o, fail entirely in action order (AO) and action count (AC) tasks. Additionally, most approaches struggle with the summarization task, which require recalling multiple nuanced details from long videos.
Finally, the existing long-videos MLLMs (except for GPT-4o) are no better than those primarily trained from short-video or image related instructions. Although such models can intake much longer videos (i.e. much more frames than other baselines), it turns out that the extended input information is not effectively utilized. On the contrary, Video-LLaVA can achieve a relatively competitive performance, especially in multiple-choice, despite that it can only intake 8 frames.
@@
Methods Input Holistic Single Detail Multi Detail M-Avg G-Avg
TR AR VS NQA ER PQA SSC AO AC
Full mark - 100 100 10 100 100 100 100 10 100 100 10
Image MLLMs
[Uncaptioned image] Otter-I [18] 16 frm 25.0 25.0 2.18 25.1 25.0 24.9 4.12 13.1 25.2 23.3 3.15
[Uncaptioned image]LLaVA-1.6 [25] 16 frm 60.6 41.0 2.11 43.1 38.4 41.0 4.35 25.5 25.7 39.3 3.23
[Uncaptioned image]InternVL-1.5 [7] 16 frm 78.8 67.0 3.16 52.7 43.5 54.4 4.88 32.8 23.8 50.4 4.02
[Uncaptioned image]Claude-3-Opus [2] 16 frm 67.2 43.5 3.11 21.6 40.2 47.8 3.66 18.2 16.7 36.5 3.39
[Uncaptioned image]Qwen-VL-Max [4] 16 frm 67.4 63.5 2.71 40.3 40.9 43.3 5.21 25.0 14.8 42.2 3.96
[Uncaptioned image] GPT-4 Turbo [1] 16 frm 79.5 68.0 4.49 45.9 47.4 60.6 6.21 26.5 16.1 49.2 5.35
Short Video MLLMs
[Uncaptioned image]Otter-V [18] 16 frm 24.6 26.0 2.38 28.2 27.6 22.3 4.23 15.1 26.7 24.4 3.31
[Uncaptioned image]mPLUG-Owl-V [49] 16 frm 28.0 25.0 2.36 24.5 31.8 27.3 5.31 21.2 23.3 25.9 3.84
[Uncaptioned image]VideoChat [20] 16 frm 33.0 32.0 2.31 27.0 32.1 27.6 5.01 24.3 28.6 29.2 3.66
[Uncaptioned image]Video-ChatGPT [29] 100 frm 26.9 24.0 2.31 40.3 42.0 29.9 5.48 25.1 31.1 31.3 3.90
[Uncaptioned image]Video-LLaMA-2 [53] 16 frm 54.5 41.5 2.34 39.4 33.5 35.4 5.22 18.5 25.7 35.5 3.78
[Uncaptioned image]VideoChat2 [21] 16 frm 74.6 51.5 2.57 42.0 47.4 43.8 5.04 22.8 29.6 44.5 3.81
[Uncaptioned image]Video-LLaVA [23] 8 frm 71.6 57.0 2.43 53.2 45.2 48.4 5.25 20.1 35.9 47.3 3.84
Long Video MLLMs
[Uncaptioned image]MovieChat [36] 2048 frm 29.5 25.0 2.33 24.2 24.7 25.8 3.23 28.6 22.8 25.8 2.78
[Uncaptioned image]Movie-LLM [37] 1 fps 30.0 29.0 2.88 29.6 24.7 24.1 5.00 20.5 24.8 26.1 3.94
[Uncaptioned image]TimeChat [35] 96 frm 23.1 27.0 2.54 24.5 28.4 25.8 4.29 24.7 32.0 30.9 3.42
[Uncaptioned image]LLaMA-VID [22] 1 fps 50.8 34.5 3.22 30.1 32.7 32.5 5.22 23.9 27.8 33.2 4.22
[Uncaptioned image]MA-LMM [12] 1000 frm 51.9 35.5 2.12 43.1 38.9 35.8 4.80 25.1 24.3 36.4 3.46
[Uncaptioned image] MiniGPT4-Video [3] 90 frm 70.9 52.5 2.64 49.0 48.6 44.5 4.07 23.2 23.0 44.5 3.36
[Uncaptioned image]GPT-4o [33] 0.5 fps 87.4 74.5 4.90 64.8 57.1 65.1 6.69 56.7 46.3 64.6 5.80
[…]
Figure 3: Experimental performances on MLVU under various video lengths: GPT4-turbo (orange), MiniGPT4-Video (red), VideoChat2 (green), Video-LLaMA-2 (blue), and MA-LMM (cyan).
4.3 Detailed Analysis
We analyze the impact from video length and three factors: context length, image understanding (IU) ability, LLM-backbone. All the factors are empirically critical to MLLMs’ LVU performances.
In the first place, we evaluate MLLMs’ performances across various video lengths. For this purpose, we introduce a derivative dataset alongside MLVU, called MLVU Time-ladder. In this dataset, the same kinds of evaluation tasks are created for videos of variant lengths, including 180s, 360s, and 600s (more details presented in Appendix D). As shown in Figure 3, the performances of all models tend to decline as the video length grows, which indicates that the existing MLLMs’ LVU abilities are severely constrained by the video length. By comparison, image models, like GPT-4-turbo [1], and short-video models, like VideoChat2 [21], VideoLLaMA [53], are more vulnerable to the growth of video length, while long video models, like MiniGPT4-Video, can be relatively more resilient.
@@ 4.3 Detailed Analysis
  MGV 16 41.1 Otter 32.2 24.4 VLM2 Vicuna-7B 24.3
90 44.5$\uparrow$3.4 LLaVA 35.8 39.3$\uparrow$14.9 Vicuna-13B 35.5$\uparrow$11.2
GPT-4o 16 54.2 GPT4V 58.1 49.2 MGV LLAMA-7B 33.7
256 64.6$\uparrow$10.4 GPT-4o 63.8 54.2$\uparrow$5.0 Mistral-7B 44.5 $\uparrow$10.8
Table 2: Detailed analysis about the impact from context length, image understanding (IU) ability, and LLM Backbone. MGV: MiniGPT4-Video, GPT4V: GPT-4-turbo, VLM2: Video-LLaMA-2.
We further examine MLLMs’ performances (measured by M-Avg) under varying context lengths. Specifically, we increase MiniGPT4-Video’s input from 16 to 90 frames and GPT-4o’s input from 16 to 256 frames (the left side of Table 2). As the input length extends, both models consistently show improved performance. To investigate the impact of MLLMs’ image understanding (IU) ability, we reference the experiment results from MMMU [51] (the middle of Table 2). It’s evident that the MLLMs’ LVU performances basically align with their image understanding performances in MMMU. Finally, we compare MLLMs with different backbones (the right side of Table 2). The results show that LVU performances improve with larger (Vicuna-13B vs. Vicuna-7B) and more capable (Mistral-7B vs. Llama-2-7B) backbone encoders. These observations indicate that LVU is the result of multiple complex factors, with the ability to perceive longer videos and effectively utilize the perceived information being crucial for the improvement of LVU.
5 Conclusion & Discussion

### v3

@@ 3.2.2 Single-Detail LVU
Needle Question-Answering (NQA). Needle-In-the-Haystack-Search (NIHS) is a popular evaluation task for long-context LLM [31]. Taking the inspiration from NIHS, we create Needle Question-Answering (NQA), shown as Figure 2 (d). In this task, the MLLM is required to answer a question related to a specific segment (referred as needle) within a long video (referred as background video). The needles are short video clips sampled from WebVid [5] and Clevrer [55], while the background videos are sampled from our ULVC. The needle is randomly inserted into the background video, where a question-answer pair is annotated. By incorporating necessary details, the question can always correspond to the needle without ambiguity. During evaluation, the MLLM needs to infer the location of the needle based on the details provided in the question, and solve the problem on top of the needle’s information. The NQA task is structured as multiple-choice, whose performance is measured by accuracy.
Ego Reasoning (ER). Ego-centric videos capture a series of consecutive actions from a first-person perspective. The MLLM needs to reason for a question about a specific behavior in the video, e.g., predicting for the event which is correlated or satisfies a certain causal relationship with the behavior (Figure 2 e). Both videos and QA annotations are collected from the NLQ task of Ego4D [15]. The ER task is structured as multiple-choice, with a total of 405 questions created for this task.
Plot Question-Answering (PQA). In this task, the MLLM needs to reason for questions about a plot in a narrative video, shown as Figure 2 (f). The video is sampled from the movies, TV series, and animated series in our ULVC. There are 589 question-answer pairs created by manual annotation. During annotation, the human annotators are asked to only provide necessary details about the plot but not to suggest any objective hints, e.g., the two characters in the example video are referred as cat and mouse, rather than Tom and Jerry. Therefore, it can prevent the question from being short-cut by the MLLM’s common-sense knowledge (more details about PQA can be found in the Appendix F.6).
Sub-Scene Captioning (SSC). In this task, the MLLM needs to generate the caption for a sub-scene in a long video. The long videos in SSC are sampled from the Movie101 dataset [58], while the questions and answers are manually annotated. During annotation, the human annotator is asked to provide a detailed description for the sub-scene as the ground-truth answer. Besides, they need to offer necessary clues in their questions such that the referred sub-scenes can be identified without ambiguity. During evaluation, we employ GPT-4 [1] to measure the quality of caption in comparison with the ground-truth. Details about annotation and evaluation are presented in Appendix F.7 and G.3.
@@ 3.2.3 Multi-Detail LVU
Action Order (AO). In this task, the MLLM needs to predict the right order for a sequence of actions (Figure 2 h). The actions are presented by short video clips, called probes. The probes are formulated in two different ways. One is made up of clips from the Kinetics dataset [18], where each clip represents a distinct action. The other one is from the consecutive clips of an action in the ActivityNet-Caption dataset [20]. The probes are inserted into a long background video, which is sampled from ULVC. There are 329 AO questions in total. The task is structured as a multiple-choice prblem, where the right order is selected from the misleading options provided by the annotator.
Action Count (AC). This task requires the MLLM to count the occurrences of an action within a long video (Figure 2 i). Each action corresponds to multiple short probe clips sampled from the Kinetics dataset [18]. The probes of an action are inserted into a long background video sampled from ULVC. We also perform manual examination to ensure that the inserted action does not exist in the original background video. A total of 266 evaluation instances have been created. The AC task is structured as a multiple-choice problem, with performance measured by accuracy.
Methods Date Input Holistic Single Detail Multi Detail M-Avg G-Avg
TR AR VS NQA ER PQA SSC AO AC
Full mark 100 100 10 100 100 100 10 100 100 100 10
Random 16.7 16.7 16.7 16.7 16.7 16.7 16.7 16.7
Image MLLMs
Otter-I [23] 2023-05 16 frm 17.6 17.9 2.03 16.7 17.0 18.0 3.90 15.7 16.7 17.1 2.97
LLaVA-1.6 [29] 2024-01 16 frm 63.7 17.9 2.00 13.3 26.4 30.0 4.20 21.4 16.7 27.1 3.10
InternVL-2 [8] 2024-07 16 frm 85.7 51.3 2.55 48.3 47.2 52.0 5.25 32.9 15.0 47.5 3.90
Claude-3-Opus [2] 2024-03 16 frm 53.8 30.8 2.83 14.0 17.0 20.0 3.67 10.0 6.7 21.8 3.25
Qwen-VL-Max [4] 2024-01 16 frm 75.8 53.8 3.00 15.0 26.4 4.84 20.0 20.7 11.7 32.2 3.92
Short Video MLLMs
Otter-V [23] 2023-05 16 frm 16.5 12.8 2.18 16.7 22.6 22.0 4.20 12.9 13.3 16.7 3.19
mPLUG-Owl-V [54] 2023-04 16 frm 25.3 15.4 2.20 6.7 13.2 22.0 5.01 14.3 20.0 16.7 3.61
VideoChat [25] 2023-05 16 frm 26.4 12.8 2.15 18.3 17.0 22.0 4.90 15.7 11.7 17.7 3.53
Video-LLaMA-2 [59] 2024-08 16 frm 52.7 12.8 2.23 13.3 17.0 12.0 4.87 15.7 8.3 18.8 3.55
VideoChat2-HD [26] 2024-06 16 frm 74.7 43.6 2.83 35.0 34.0 30.0 5.14 21.4 23.3 37.4 3.99
Video-LLaVA [28] 2023-11 8 frm 70.3 38.5 20.9 2.30 26.4 26.0 5.06 20.0 21.7 29.3 3.68
ShareGPT4Video [7] 2024-05 16 frm 73.6 25.6 2.53 31.7 45.3 38.0 4.72 17.1 8.3 34.2 3.63
VideoLLaMA2 [9] 2024-06 16 frm 80.2 53.8 2.80 36.7 54.7 54.0 5.09 42.9 16.7 48.4 3.95
Long Video MLLMs
MovieChat [41] 2023-07 2048 frm 18.7 10.3 2.30 23.3 15.1 16.0 3.24 17.1 15.0 16.5 2.77
Movie-LLM [42] 2024-03 1 fps 27.5 25.6 2.10 10.0 11.3 16.0 4.93 20.0 21.7 18.9 3.52
LLaMA-VID [27] 2023-11 1 fps 20.9 23.1 2.70 21.7 11.3 16.0 4.15 18.6 15.0 18.1 3.43
MA-LMM [16] 2024-04 1000 frm 44.0 23.1 3.04 13.3 30.2 14.0 4.61 18.6 13.3 22.4 3.83
MiniGPT4-Video [3] 2024-04 90 frm 64.9 46.2 2.50 20.0 30.2 30.0 4.27 15.7 15.0 31.7 3.39
LongVA [60] 2024-06 256 frm 81.3 41.0 2.90 46.7 39.6 46.0 4.92 17.1 23.3 42.1 3.91
Video-CCAM [11] 2024-08 96 frm 79.1 38.5 2.65 45.0 52.8 56.0 4.49 24.3 26.7 46.1 3.57
Video-XL [40] 2024-09 256 frm 78.0 28.2 3.40 50.0 41.5 46.0 5.02 48.6 31.7 46.3 4.21
LLaVA-Onevision [24] 2024-08 32 frm 83.5 56.4 3.75 46.7 58.4 58.0 5.09 35.7 23.3 51.7 4.42
GPT-4o [37] 2024-05 0.5 fps 83.7 68.8 4.94 42.9 47.8 57.1 6.80 46.2 35.0 54.5 5.87
[…]
@@ 4.1 Settings
We conduct a comprehensive investigation of 23 MLLMs using our MLVU benchmark, encompassing both open-source and proprietary models. The experimental MLLMs are divided into three categories: 1) Image MLLMs, primarily fine-tuned using image-related instructions; 2) Short Video MLLMs, fine-tuned with short-video related instructions; and 3) Long Video MLLMs, optimized for long-video understanding capability. For Image MLLMs, we leverage their multi-image inference capabilities to process segmented frames from original videos. For Video MLLMs, we employ either a uniform sampling strategy or a frame rate sampling strategy for video processing. All models are evaluated based on their official implementations or available APIs, with evaluations conducted in a zero-shot manner. More details about the evaluation are provided in Appendix G.
4.2 Main Results
@@
The overall evaluation results for all investigated MLLMs in the MLVU test set are shown in Table 2 (with dev set results in Appendix B). Individual performances are reported for each task, while average performances are provided for multiple-choice (M-Avg) and generation tasks (G-Avg). From the results, we derive three primary conclusions:
1) The proprietary model GPT-4o [37] achieves optimal performance in our benchmark. It leads in multiple-choice tasks with an M-Avg of 54.5%(within 0-100%) and excels in generation tasks with a G-Avg of 5.87 (within 0.0-10.0), outperforming all other methods.
2) Recent advances in LVU have achieved significant progress, and the gap between open-source long video MLLMs and GPT-4o on close-ended tasks is narrowing. Before June 2024, the best open-source long video MLLMs, MiniGPT4-Video [3], lagged significantly behind GPT-4o. However, recent models [60, 11, 40, 24] have made substantial progress. For instance, LLaVA-Onevision trails GPT-4o by only 2.8% in M-Avg. These models have improved their ability to handle long visual sequences, achieving significant advancements in single-detail (e.g., NQA) and multi-detail (e.g., AC) tasks compared to previous open-source models.
@@
3) Existing methods still struggle to handle most tasks in our benchmark. For instance, GPT-4o only achieves 42.9% in the needle question-answering (NQA) task. In contrast, analogous tasks in the text domain, such as NIHS (Needle-In-the-Haystack-Search) and Passkey Retrieval, are effectively handled by many existing long LLMs [14, 61]. Additionally, GPT-4o shows even less reliability in tasks like ego-reasoning (ER), action ordering (AO), and action count (AC), with most baseline methods performing even worse. These observations indicate that long-video understanding remains a significant challenge for today’s MLLMs.
In addition to the primary conclusions from the overall performances, we can also make the following interesting observations about the individual tasks.
@@
4) The close-ended holistic tasks present much higher differentiation than other tasks. These tasks, i.e., topic reasoning (TR) and anomaly recognition (AR), show significant variance in performance across different models. Proprietary MLLMs, like GPT-4o, and superior open-source models, such as InternVL-2 [8], VideoLLaMA2 [9], and LLaVA-OneVision [24], can accurately solve these problems. Meanwhile, many other popular MLLMs still fail to generate meaningful performances. Since these tasks only require an overall understanding of long videos, they can serve as a preliminary indicator of MLLMs’ long video understanding (LVU) ability.
@@
5) It’s challenging to deal with tasks that require nuanced understanding of multiple details. Although several MLLMs can handle single-detail LVU tasks to some extent, their performances suffer from catastrophic degradation when addressing multi-detail LVU tasks. Most methods, except for GPT-4o and Video-XL [40], fail entirely in action order (AO) and action count (AC) tasks. Additionally, most approaches struggle with summarization tasks, which require recalling multiple nuanced details from long videos.
@@
4.3 Further Analysis
@@
  MGV 16 24.2 Otter-I 32.2 17.1 VLM2 Vicuna-7B 13.3
90 31.7$\uparrow$7.5 LLaVA-1.6 35.8 27.1$\uparrow$10.0 Vicuna-13B 18.8$\uparrow$5.5
GPT-4o 16 45.8 GPT-4V 58.1 43.3 MGV LLaMA-7B 20.6
256 54.5$\uparrow$8.7 GPT-4o 63.8 45.8$\uparrow$2.5 Mistral-7B 31.7 $\uparrow$11.1
Table 3: Detailed discussions about the impact from context length, image understanding (IU) ability, and LLM Backbone. For the IU impact experiment, we used 16-frame uniform sampling for both GPT-4V and GPT-4o. MGV: MiniGPT4-Video, VLM2: Video-LLaMA-2.
6) Longer videos are more challenging for MLLMs. We evaluate MLLMs’ performances across various video lengths. For this purpose, we introduce a derivative dataset alongside MLVU, called MLVU Time-ladder. In this dataset, the same kinds of evaluation tasks are created for videos of variant lengths, including 180s, 360s, and 600s (more details presented in Appendix D). As shown in Figure 3, the performances of all models tend to decline as the video length grows, which indicates that the existing MLLMs’ LVU abilities are severely constrained by the video length. Moreover, the short video model Video-LLaMA-2 [59] maintains a certain level of LVU ability at 3 minutes, but its performance approaches random results at 10 minutes.
7) The performance of recent advanced long video MLLMs remains robust regardless of the position of the referring clip within the long video. In single-detail tasks, the referring clip denotes the specific segment of the long video that is referenced or inferred to answer a question. As shown in Figure 4, we categorize clip positions into four intervals and assess model performance on two single-detail tasks: ego reasoning (ER) and plot question-answering (PQA). Recent long video MLLMs, such as LongVA [60] and Video-XL [40], maintain consistent performance regardless of the referring clip’s position within the video. Conversely, short video MLLMs are more sensitive to clip location. This indicates that recent advancements in long video MLLMs enhance both reliable clue retrieval and effective reasoning from extended visual sequences.
Figure 3: Experimental performance on varying video lengths. The evaluated metric is the average accuracy across five multiple-choice tasks involving local information: NQA, ER, PQA, AC, and AO.
Figure 4: Model performance across different referring clip positions, spanning from the beginning to the end of the entire video.
8) The challenge of multi-detail tasks increases with the number of details. We analyzed model performance on the action count (AC) task by grouping questions based on the number of probes (which correspond to details) and evaluating the average performance within these groups. As shown in Figure 5, performance significantly declines across all models as the number of probes increases. This indicates that current MLLMs face substantial difficulties comprehending and processing multiple details simultaneously, highlighting a critical area for future improvement in long video understanding capabilities.
Figure 5: Model performance on the action count (AC) task in relation to the number of probes.
9) Context Length, Image-Understanding ability, and the choice of LLM Backbones are key factors in LVU performance. As shown in Table 3, we conducted ablation experiments on several factors affecting MLLMs, using M-Avg as the evaluation metric. First, we examined the models’ handling of different context lengths. Specifically, we increased MiniGPT4-Video’s input from 16 to 90 frames and GPT-4o’s input from 16 to 256 frames (as shown on the left side of Table 3). Both models showed consistent performance improvements with longer input lengths. To assess the impact of image understanding (IU) capabilities, we referred to the results from MMMU [57] (presented in the middle of Table 3). It is evident that MLLMs’ LVU performance generally aligns with their IU performance in MMMU. Finally, we compared MLLMs using different backbones (depicted on the right side of Table 3). The findings indicate that LVU performance improves with larger (Vicuna-13B vs. Vicuna-7B) and more advanced backbones (Mistral-7B vs. Llama-2-7B). These observations indicate that LVU is the result of multiple complex factors, with the ability to perceive longer videos and effectively utilize the perceived information being crucial for the improvement of LVU.
5 Conclusion